### v1

# Andrés Vallejo    Alejandro Romanelli    Raúl Donangelo Facultad de Ingeniería, Universidad de la República, Montevideo, Uruguay

Abstract
We derive a generalization of the Ehrenfest theorem valid for open quantum systems. From this result, we identify three factors contributing to the evolution of expected values: explicit time dependence of the observable, thermal interaction, and quantum coherence. When considering the local Hamiltonian as the observable, we obtain an alternative version of the first law of thermodynamics. In some cases, the non-thermal contributions to the energy rate of change can be expressed as the expected value of a Hermitian operator, so the power performed by the system can be considered a quantum observable. As an application, the pure dephasing process is reinterpreted from this perspective.
Keywords: Ehrenfest Theorem, quantum thermodynamics, mechanical power, observables.
Corresponding author: Andrés Vallejo
E-mail address: avallejo@fing.edu.uy
I Introduction
A well-known result in quantum theory states that, for a system in the state $|\psi\rangle$, the time derivative of the expected value of any observable quantity $O$ is given by the expression ($\hbar=1$) [1, 2]:
$\dfrac{d\langle O\rangle}{dt}=\left\langle\dfrac{\partial O}{\partial t}\right\rangle+i\left\langle[H,O]\right\rangle,$ (1)
where $\langle\bullet\rangle=\langle\psi|\bullet|\psi\rangle$ represents the expected value of the operator $\bullet$, $H$ is the Hamiltonian of the system, and $[H,X]$ is the commutator between $H$ and $X$. This result has important applications. For instance, when applied to the position and momentum operators $X$ and $P$ of a particle with mass $m$, subjected to the potential $V(X)$, it leads to the famous Ehrenfest’s Theorem [3]:
$\dfrac{d\langle X\rangle}{dt}=\dfrac{\langle P\rangle}{m},\hskip 8.5359pt\dfrac{d\langle P\rangle}{dt}=-\left\langle\dfrac{\partial V}{\partial X}\right\rangle,$ (2)
which, in some cases, provides support to the correspondence principle [4]. Eq. (1) is also the starting point for the derivation of important results involving the quantum speed limit [5].
Another obvious but relevant consequence of Eq. (1) is that any time-independent observable that commutes with the Hamiltonian of a closed system is a constant of motion. In particular, the internal energy, defined as the expected value of the Hamiltonian,
$E=\langle H\rangle$ (3)
is a conserved quantity, provided that $H$ does not depend explicitly on time. This is consistent with the fact that closed classical systems have constant energy. However, when open systems are considered, the energy of the system is no longer a conserved quantity due to the exchange of heat and work with the environment. In that case, it would be interesting to evaluate the effects of these interactions in the dynamics of expected values. In what follows, we derive a generalization of Eq. (1) valid for an arbitrary open dynamics that allows us to address this issue.
The outline of this paper is as follows. In Section II we obtain the generalization of Eq. (1) for the case of a generic open dynamics. From this result, we identify the mechanical, thermal and coherence-related contributions to the change of observable quantities of the system. In Section III, we show how the application of this result to the particular case in which the observable is the Hamiltonian of the system, leads to an alternative version of the first law of thermodynamics. We then focus on the study of two-level systems and show that the power of coherence can be interpreted as rotational work per unit of time in the presence of an external field. We also discuss the dephasing process from this perspective. Remarks and conclusions are presented in Section IV.
II Time evolution of expected values in an arbitrary open dynamics
II.1 Derivation
The statistical properties of an open system are described by its reduced density operator $\rho$, in terms of which the expected value of any observable quantity $O$ can be written as:
$\langle O\rangle=\text{tr}\left(\rho O\right),$ (4)
where tr denotes the trace operation. Let us consider the spectral decomposition of $\rho$,
$\rho=\sum_{j}\lambda_{j}|\psi_{j}\rangle\langle\psi_{j}|,$ (5)
where $|\psi_{j}\rangle$ is the eigenvector of $\rho$ corresponding to the eigenvalue $\lambda_{j}$. From Eqs. (4) and (5), we obtain that
$\langle O\rangle=\sum_{j}\lambda_{j}\langle\psi_{j}|O|\psi_{j}\rangle,$ (6)
so an infinitesimal change in the expected value can be written as
$d\langle O\rangle=\sum_{j}(d\lambda_{j})\langle\psi_{j}|O|\psi_{j}\rangle+\sum_{j}\lambda_{j}d(\langle\psi_{j}|O|\psi_{j}\rangle).$ (7)
Given that $\rho$ is Hermitian, its set of eigenvectors evolves in time preserving the orthonormal-basis structure. Consequently, there exists a unitary operator $T$, possibly time-dependent, such that, for each eigenvector of $\rho$, satisfies that
$|\psi_{j}(t+dt)\rangle=T(t,t+dt)|\psi_{j}(t)\rangle.$ (8)
Since $d$t is small, $T$ is a unitary operator close to the identity, so we can approximate it by [6]:
$T(t,t+dt)=\mathbb{I}-i\Omega(t)dt,$ (9)
where $\mathbb{I}$ is the identity operator and $\Omega$ is the Hermitian generator of the eigenbasis’s dynamics. Using Eqs. (8) and (9), we obtain that
$\begin{cases}d|\psi_{j}\rangle=|\psi_{j}(t+dt)\rangle-|\psi_{j}(t)\rangle=-i\Omega|\psi_{j}\rangle dt\\ d\langle\psi_{j}|=\langle\psi_{j}(t+dt)|-\langle\psi_{j}(t)|=i\langle\psi_{j}|\Omega dt,\end{cases}$ (10)
so from Eqs. (7) and (10), the variation in the expected value of $O$ can be written as
$\begin{split}d\langle O\rangle&=\sum_{j}d\lambda_{j}\langle\psi_{j}|O|\psi_{j}\rangle+\sum_{j}\lambda_{j}\langle\psi_{j}|dO|\psi_{j}\rangle\\ &+i\sum_{j}\lambda_{j}\langle\psi_{j}|[\Omega,O]|\psi_{j}\rangle dt.\end{split}$ (11)
Therefore the time-derivative of $\langle O\rangle$ adopts the form
$\begin{split}\dfrac{d\langle O\rangle}{dt}&=\sum_{j}\dot{\lambda_{j}}\langle\psi_{j}|O|\psi_{j}\rangle+\left\langle\dfrac{\partial O}{\partial t}\right\rangle+i\left\langle[\Omega,O]\right\rangle,\end{split}$ (12)
which is a natural generalization of Eq. (1) for open systems.
II.2 Discussion
When comparing Eqs. (1) and (12), we note two important differences. First, the appearance of the term
$\sum_{j}\dot{\lambda_{j}}\langle\psi_{j}|O|\psi_{j}\rangle,$ (13)
which involve the time derivative of the eigenvalues of $\rho$. Since the evolution of a closed system is unitary, and noting that unitarity preserves the natural populations, we conclude that this contribution is necessarily related to some kind of interaction with the environment. One possible interpretation of this term arises recalling that the von Neumann entropy of the system can be written as
$S_{vN}=-k_{B}\sum_{j}\lambda_{j}\log\lambda_{j},$ (14)
where $k_{B}$ is the Boltzmann constant, so its rate of change adopts the form
$\dot{S}_{vN}=-k_{B}\sum_{j}\dot{\lambda}_{j}\log\lambda_{j}.$ (15)
Eqs. (13) and (15) are clearly connected (albeit in an intricate way), since both are linear combinations of the set $\{\dot{\lambda}_{j}\}$. Hence, assuming that the von Neumann entropy is a valid extension of the thermodynamic entropy, and considering that the contribution given by Eq. (13) is only present when there is an entropy variation, we conclude that it could be considered a contribution of thermodynamic nature, linked to the heat exchanged between the system and the environment. However, we must note that the identification of von Neumann entropy with the thermodynamic entropy is only confirmed for Gibbs states, so the extension of this identification to other states has been controversial [18, 19, 20].
On the other hand, the remaining terms in Eq. (12):
$\left\langle\frac{\partial O}{\partial t}\right\rangle+i\left\langle[\Omega,O]\right\rangle$ (16)
correspond to those appearing in Eq. (1), with $\Omega$ playing the role of $H$. This is so because it is $\Omega$, instead of $H$, that governs the evolution of the eigenvectors of $\rho$ in an open dynamics. The first part,
$\left\langle\dfrac{\partial O}{\partial t}\right\rangle$ (17)
is associated with the possible explicit time dependence of the observable, for example due to experimental control. The second part is discussed in detail in the following.
The contribution of coherence
In order to interpret the term
$i\langle[\Omega,O]\rangle=i\text{tr}(\rho[\Omega,O]),$ (18)
we first focus the simple case of a two-level system. In that case, the density operator can be written in terms of the Bloch vector, $\vec{B}=\langle\vec{\sigma}\rangle$, as
$\rho=\dfrac{1}{2}(\mathbb{I}+\vec{B}.\vec{\sigma}),$ (19)
where $\vec{\sigma}$ is a formal vector whose components are the Pauli matrices, and $\mathbb{I}$ is the identity operator. Similarly, the operators $\Omega$ and $\vec{O}$ can be written, aside from irrelevant terms which are multiples of the identity, as:
$\Omega=-\vec{w}.\vec{\sigma},$ (20)
and
$O=-\vec{o}.\vec{\sigma}$ (21)
This representation of 2-dimensional operators by vectors in $R^{3}$ has many advantages. For example, it is easy to verify that two observables commute if and only if their representative vectors are collinear. Using the identity
$(\vec{a}.\vec{\sigma})(\vec{b}.\vec{\sigma})=(\vec{a}.\vec{b})\mathbb{I}+i(\vec{a}\times\vec{b}).\vec{\sigma},$ (22)
we obtain that
$[\Omega,O]=2i(\vec{w}\times\vec{o}).\vec{\sigma},$ (23)
and finally, from Eqs. (19), (22) and (23), it can be shown that
$i\langle[\Omega,O]\rangle=2\vec{B}.(\vec{o}\times\vec{w}).$ (24)
so the third contribution to the evolution of expected values is proportional to the triple product of the vectors $\vec{B}$, $\vec{o}$ and $\vec{w}$.
Since the triple product can be geometrically interpreted as the volume of the parallelepiped defined by the three vectors, it is clear that this contribution is present only when the set formed by these vectors is linearly independent. In particular, this implies that the density operator must necessarily have coherence in both the $\Omega$ and $O$ eigenbasis. However, the presence of coherence is not sufficient, since the volume can be zero provided that $\vec{B}$ lies in the plane defined by the other two vectors. On the other hand, the maximum contribution to the rate of change of $\langle O\rangle$ due to the presence of coherence occurs when the three vectors $\vec{o}$, $\vec{w}$ and $\vec{B}$ define a rectangular box, and its value is:
$|i\langle[\Omega,O]\rangle|_{max}=2|\vec{w}||\vec{o}||\vec{B}|$ (25)
In $N$ dimensions, the cyclic property of the trace implies that
$\text{tr}(\rho[\Omega,O])=\text{tr}(O[\rho,\Omega])=\text{tr}(\Omega[O,\rho]),$ (26)
so again we find that, for this contribution to be present, i) $\rho$ cannot commute with $O$ and $\Omega$ (it must have coherence in the eigenbases of both operators); ii) the set formed by these three operators must be linearly independent. We conclude that this contribution is a measure of some kind of set coherence associated with the observables $\rho$, $\Omega$ and $O$.
Finally, it is also interesting to notice that, for time-independent observables, the application of the generalized uncertainty principle [7],
$\sigma_{\Omega}\sigma_{O}\geq\dfrac{1}{2}|\left\langle[\Omega,O]\right\rangle|,$ (27)
establishes that the upper bound of $|\left\langle[\Omega,O]\right\rangle|$ is proportional to the product of the corresponding variances, when a measurement of both operators is performed:
$|\left\langle[\Omega,O]\right\rangle|\leq 2\sigma_{\Omega}\sigma_{O}.$ (28)
III Energetic Implications
Applying the previous formalism to the particular case in which the observable $O$ is the local Hamiltonian of the system, and defining the internal energy as the expected value of the Hamiltonian, $E=\langle H\rangle$, we obtain that
$\dfrac{dE}{dt}=\sum_{j}\dot{\lambda_{j}}\langle\psi_{j}|H|\psi_{j}\rangle++\left\langle\mathbb{P}_{H}\right\rangle,$ (29)
where $\mathbb{P}_{H}$ is the Hermitian operator:
$\mathbb{P}_{H}=\dfrac{\partial H}{\partial t}+i[\Omega,H].$ (30)
In light of the discussion carried out in the previous section, it is clear that the first term on the right-hand side of Eq. (29) corresponds to the rate of change of the internal energy that is accompanied by an entropy variation, so we interpret this term as the heat flux per unit of time exchanged by the system:
$\dot{\mathcal{Q}}\equiv\sum_{j}\dot{\lambda}_{j}\langle\psi_{j}|H|\psi_{j}\rangle,$ (31)
as it was recently proposed in Refs. [10, 11, 9].
Consequently, assuming that the classical statement of the first law
$\frac{dE}{dt}=\dot{\mathcal{Q}}+\dot{\mathcal{W}}$ (32)
still holds, we conclude that the second term in Eq. (29) is the total power performed on the system:
$\dot{\mathcal{W}}\equiv\langle\mathbb{P}_{H}\rangle.$ (33)
The power includes the standard mechanical contribution $\langle\partial H/\partial t\rangle$ due to Hamiltonian driving, plus the coherence power $i\langle[\Omega,H]\rangle$, which is related to the level of incompatibility between the local Hamiltonian $H$, which governs the free evolution of the system, and the effective Hamiltonian $\Omega$ that governs the evolution of the eigenbasis of $\rho$ when the interactions with the environment are included. Accordingly, the coherence power is related to the work needed to deviate the quantum trajectory from the natural evolution defined by $H$.
Naturally, if the rotation of the eigenstates of $\rho$ is not modified by the interaction with the environment, $\Omega$ and $H$ commute. Therefore, in this case, no additional power is required to produce the new evolution. However, in the general case, both evolutions will be different, and since the local Hamiltonian $H$ fixes the natural evolution of the state of the system (and therefore, of the eigenbasis), deviations from that behaviour may imply an energy exchange with the environment, which can be evaluated as the expected value of the commutator between both operators.
Finally, it is important to notice that, in many situations of interest, both $H$ and $\Omega$ are time-independent. In those cases, the total power corresponds to the expected value of the time-independent Hermitian operator $\mathbb{P}_{H}$, so it can be considered as a quantum observable. For this reason we will call $\mathbb{P}_{H}$ the power operator.
III.1 Example: a two-level system
III.1.1 Physical interpretation of coherence work
Let us consider an open two-level system whose local Hamiltonian
$H=-\vec{v}.\vec{\sigma}$ (34)
is time-independent. From Eqs. (19), (20), (30), (33), and (34), and employing the cyclic property of the triple product, it is possible to show that
$\dot{\mathcal{W}}=i\left\langle[\Omega,H]\right\rangle=2\vec{w}.(\vec{B}\times\vec{v}).$ (35)
To interpret Eq. (35) physically, notice that the Hamiltonian $H$ in Eq. (34) is equivalent to that describing the evolution of a spin in a magnetic field proportional to $\vec{v}$. As a consequence, the product $\vec{B}\times\vec{v}$ represents the torque performed by the magnetic field $\vec{v}$ on the spin $\vec{B}$:
$\vec{\tau}=\vec{B}\times\vec{v}.$ (36)
Let us now focus on the other factor appearing in Eq. (35), $2\vec{w}$. Since $\vec{w}$ determines via Eq. (20) the effective Hamiltonian $\Omega$ that governs the evolution of the eigenbasis of $\rho$ in Hilbert space, it is clear that $\vec{w}$ should be related to the evolution of the state in the Bloch vector representation. To clarify this point, let us note that the evolution of the normalized Bloch vector can be understood as an instantaneous rotation around certain vector $\vec{\Omega}$, which plays the role of an angular velocity:
$\frac{d\hat{B}}{dt}=\vec{\Omega}\times\hat{B}.$ (37)
From Ec. (34) we can write the internal energy of the system as
$E=\langle H\rangle=-\vec{B}.\vec{v},$ (38)
so the rate of change of the internal energy can be separated in two terms:
$\dfrac{dE}{dt}=-\dot{B}\hat{B}.\vec{v}-B\dfrac{d\hat{B}}{dt}.\vec{v}$ (39)
Taking into account that the entropy of a two-level system depends only on the modulus of the Bloch vector, it was shown in Ref. [14] that the first term in the right-hand side of the equation above is the heat exchanged with the environment per unit of time:
$\dot{\mathcal{Q}}=-\dot{B}(\hat{B}.\vec{v}),$ (40)
Consequently, the second term represents the power performed by the system:
$\dot{\mathcal{W}}=-B\dfrac{d\hat{B}}{dt}.\vec{v},$ (41)
which, using Eq. (36) and (37), can be written as:
$\dot{\mathcal{W}}=-\vec{\Omega}.\vec{\tau},$ (42)
so it adopts the classical form of the power needed to rotate a spin in the presence of an external field. Comparing Eqs. (35) and (42), we conclude that $-2\vec{w}$ can be identified as the angular velocity $\vec{\Omega}$.
This simple model allows to clarify some aspects previously discussed in a more general way. First, notice that if the free angular velocity of the Bloch vector, which is proportional to $\vec{v}$, and the actual angular velocity, which is proportional to $\vec{w}$ are parallel, i.e. if $H$ and $\Omega$ commute, the vectors $\vec{\tau}=\vec{B}\times\vec{v}$ and $\vec{\Omega}$ are orthogonal, so no external power is required to produce the evolution, see Fig. (1). On the other hand, it is clear that the maximum power occurs when $\vec{\tau}$ and $\vec{\Omega}$ are antiparallel. As expected, note that this contribution only exists if the system presents coherence in the energy eigenbasis, since for incoherent states the vectors $\vec{v}$ and $\vec{B}$ are parallel, and the torque given by equation (36) is zero. This result evidences the thermodynamic value of quantum coherence.
Figure 1: Geometrical interpretation of the coherence power for a two-level system. In the natural unitary evolution (left), the Bloch vector rotates around $\vec{\Omega}$, which is parallel to $\vec{v}$, so the vectors $\vec{\tau}$ and $\vec{\Omega}$ are orthogonal, and no power is required to maintain the rotation. If the vectors $\vec{\tau}$ and $\vec{\Omega}$ are not orthogonal (right), coherence power is needed to deflect the trajectory from its natural evolution.
III.1.2 Alternative analysis of a pure-dephasing process
As an example, let us analyze a pure-dephasing process. An outline of this discussion was presented in Ref. [14], and is now extended.
The pure-dephasing process, also called phase-damping, is a quantum channel commonly employed as a schematic description of the decoherence processes. In this process it is assumed that the coherence loss of an initial quantum superposition occurs due to interactions with the environment.
This environment is supposed to be composed of small subsystems, e.g. photons from some background radiation, which are scattered by the system at a certain rate $\Gamma$. If we assume that the time scale during which the energy and momentum transfer is significant is longer than $\Gamma^{-1}$, it can be shown that the process is characterized by the preservation of the populations of the system, in parallel with an exponential decay of the coherences of its density matrix, when it is expresses in the energy eigenbasis [15].
Let us consider the pure-dephasing of a two-level atom with local Hamiltonian $H=-\varepsilon\sigma_{z}$, corresponding to an effective field $\vec{v}=\varepsilon\hat{z}$. If the initial state is defined by the Bloch vector $\vec{B}(t=0)=(B_{x},0,B_{z})$, its evolution is given by [16]
$\vec{B}(t)=\left(B_{x}\cos(\Lambda t)e^{-\Gamma t},B_{x}\sin(\Lambda t)e^{-\Gamma t},B_{z}\right),$ (43)
where $\Lambda=\sqrt{2\varepsilon/\hbar}$. From Eq. (43) we notice that the projection of the Bloch vector on the $\hat{z}$ direction is constant, so the energy of the atom does not change. Considering that the Hamiltonian of the system is constant, from the standard point of view there is no work involved in the process [17]. As a consequence, due to energy conservation, there is no heat exchanged either, so the information about the phase is lost without the intermediation of heat or work.
However, the entropy increase occurring in the pure-dephasing process should imply that heat is absorbed from the environment. Integrating Eq. (40) along the trajectory given by Eq. (43), we obtain:
[…]
which, in fact, is a positive quantity for positive-temperature states ($B_{z}>0$). Since the energy is constant, the system does an equivalent amount of coherence work on the environment, which should be associated to the scattering process.
IV Final Remarks and conclusions
Since unitary evolutions preserve the eigenvalues of the density operator, the explicit time-dependence of these quantities is an accurate signature of the interaction of the system with its environment. In this work he have assumed that this interaction consists of a thermal contribution, which is associated with the entropy variation, and a power contribution that does not change the entropy of the system. Noting that the entropy depends exclusively on those eigenvalues, we were able to determine to what extent these interactions affect the values of the other properties of the system, generalizing Ehrenfest’s theorem, which is valid for unitary dynamics.
The above partition is based on the hypothesis that the von Neumann entropy is a proper extension of the thermodynamic entropy in the quantum regime. While this is strictly true for the particular case of Gibbs states, there is controversy about the validity of this identification for more general states [18, 19, 20]. We believe that, by exploring the consequences of having assumed this hypothesis, we can contribute to shed light on this issue. However, note that the validity of Eq. (12) is independent on this hypothesis, which only plays a role in the interpretation of the different terms appearing in that equation.
In addition to the possible time-dependence of observables through variable parameters, we have shown that there exists a coherence contribution linked to the level of incompatibility between the observables and a certain effective Hamiltonian that governs the evolution of the eigenvectors of the density matrix. For the particular case in which the observable considered is the Hamiltonian of the system, the contributions just described give rise to alternative versions of heat and work.
Particularly interesting is the fact that for time-independent Hamiltonians, if the interaction is such that there is also no time dependence in the effective Hamiltonian $\Omega$, the power can be calculated as the expected value of a time independent Hermitian operator, so it can be considered as a quantum observable. Additionally, for two-level systems we have shown that the coherence power is related to the deviation of the Bloch vector motion with respect to its natural tendency to rotate around the privileged direction determined by the external field. We emphasize that several very recent works employ these definitions, but derived in the different context of shortcuts to adiabacity instead of as particular case of a generalized Ehrenfest formula [10]. For example, in Ref. [21], entropy-based definitions of heat and work which are equivalent to Eqs. (40) and (42) are employed in order to quantify the non-Markovianity of quantum dynamical maps.
This interpretation allows to analyze well-known processes, such as thermalization, or pure dephasing, from an alternative point of view.
Acknowledgments
This work was partially supported by Comisión Académica de Posgrado (CAP), Agencia Nacional de Investigación e Innovación (ANII) and Programa de Desarrollo de las Ciencias Básicas (PEDECIBA) (Uruguay).